A battery of e.m.f 12V supplies a current of 5A for 2 minutes. How much energy is supplied in this time?
1
Expert's answer
2020-04-17T10:17:50-0400

From this formula e.m.f.


"\epsilon=\frac{A}{q}"

we have


"A=q\cdot\epsilon=It\epsilon"

where


"I=5 (A); t=120 (s); \epsilon=12(V)""A=5\cdot120\cdot12=7200 (J)"